Assume you run a company that offers a product for which all consumers have an identical demand curve. Each consumer’s demand curve is P = 20 – 4Q. The marginal cost of production is $4. Devise an optimal two-part tariff pricing policy.
Answer #1

ANSWER:-

Given date

Demand curve P = 20 - 4Q

Marginal cost MC= $4

Optimal two tariff pricing per unit fee

=> MC = $4

Membership fee = Consumer surplus

=1/2 * (p Q=0 - p P=MC ) Q P=MC

Substitute P values and Q values above equation

= 1/2 * (20 - 4) (4)

=1/2 * (16) (4)

=1/2 * 64

=64/2

=32

=$ 32

Therefore

units per fee    = $ 4

membership fee = $ 32

  

----------------------------------------------------------------

IF YOU WANT ANY EXPLANATIONS ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU PLEASE DO NOT GIVE DIRECT THUMBS DOWN,IF YOU SATISFIED MY WORK GIVE ME THUMBS UP RATING

*****************THANK YOU***************